OpenStudy (anonymous):

Write the equation for the line tangent to the circle x2 + y2 = 29 at the point (2, 5).

OpenStudy (anonymous):

Can someone help me?

OpenStudy (campbell_st):

ok... rewrite the equation of the circle with y as the subject... leave you answer in index form. doing this will allow you to find the 1st derivative.

OpenStudy (anonymous):

would it be y-2=-2/5x-5

OpenStudy (campbell_st):

nope how about \[y = (29 - x^2)^{\frac{1}{2}}\] so can you find the 1st derivative of the equation..?

OpenStudy (anonymous):

I dont understand it at all then.

OpenStudy (campbell_st):

ok... so are you doing calculus at the moment...?

OpenStudy (anonymous):

nope algerbra 2

OpenStudy (campbell_st):

ok... well for me... this is a calculus question, have you done any calculus...?

OpenStudy (anonymous):

nope not at all do you want me to give you the choices i was told?

OpenStudy (campbell_st):

yep go on...

OpenStudy (anonymous):

y-5=-2/5x+2 y-2=-2/5x+1 y-5=-2/5(x-2)

OpenStudy (anonymous):

and then the one i posted earlier

OpenStudy (campbell_st):

its the last choice.... I hope they give you solutions.... so you can see how it was done...

OpenStudy (anonymous):

so the bottom one of the three i posted together?

OpenStudy (campbell_st):

yep.... thats the solution.

OpenStudy (anonymous):

Thank You.